Hi, can you help me with this question as i've been figuring out for long and still could not get the answer:
In the coming school holiday,the probability for Dave to visit his brother in Ireland is 3/5 and the probability of him going on island vacation is 2/3. Calculate the probability that Dave
a) will not visit his brother nor on an island vacation
b) will only do one activity.
Pls help me, i'm doing my o levels next week thanks

Answer by stanbon(48568)

In the coming school holiday,the probability for Dave to visit his brother in Ireland (A)is 3/5
and the probability of him going on island vacation (B) is 2/3.
Calculate the probability that Dave
------------
a) will not visit his brother nor on an island vacation
P(A' and B') = P(A')*P(B') = (2/5)*(1/3) = 2/15
-------------------
b) will only do one activity.
P(A and B) = P(A)*P(B) = (3/5)(2/3)= 2/5
P(A' and B') = 2/15 (from part a above)
Therefore: P(A or B exclusively) = 1-[2/5 + 2/15] = 1-(40/75)= 1-8/15= 7/15
===============
Cheers,
Stan H.